# Success Probability in Shor’s Algorithm

### v1

@@ 1 Introduction
Shor’s algorithm, marked a turning point in quantum computing by offering an exponentially faster method for integer factorization [20], requires the use of a pseudo-random integer satisfying three specific properties. The algorithm consists of three steps. A lower bound on the success probability of the algorithm is given in [19], but no exact value nor a closed form formula.
The goal of this paper is to compute the exact success probability for each of the three steps of Shor’s algorithm, as well as for the full algorithm, across all integers — including those that are not simply of the form $pq$. We provide the probability values for all possible cases, based on the prime factor decomposition of the input integer. We also identify all specific failure cases occurring in the different steps of Shor’s algorithm.

### v2

@@ 1 Introduction
Shor’s algorithm marked a turning point in quantum computing by offering an exponentially faster method for integer factorization [22]. The algorithm relies on the selection of a pseudo-random integer satisfying certain properties and proceeds in three main steps: random selection, quantum order finding, and classical post-processing. A lower bound on the success probability was given in [21], but no closed-form expression or exact value is known.
The goal of this paper is to compute the exact success probability for each of the three steps of Shor’s algorithm, as well as for the full algorithm, across all integers—including those not simply of the form $N=pq$. We provide these values in closed form for all relevant cases based on the prime factor decomposition of the input. Furthermore, we identify and classify all specific failure cases that may occur in different stages of the algorithm.
We emphasize that our analysis is based on Shor’s original algorithm. However, more efficient classical post-processing methods have since been developed that allow complete factorization from a single order-finding call with high probability [5, 6]. Implementations of these improved techniques are available in open-source libraries such as Quaspy and Factoritall.